# Global stellarator coil optimization with quadratic constraints and objectives

### v4

@@ 2.2 QUADCOIL
$\begin{split}&\min_{\Phi^{\prime},a,b_{j}}\left[f(\Phi^{\prime},a)+\sum_{j}^{n_{\text{cons}}}c_{j}b_{j}\right]\\ &\text{subject to }g_{j}(\Phi^{\prime},a)+d_{j}b_{j}\leq 0,\\ &\text{for all }j=1,...,n_{\text{cons}}.\end{split}$ (8)
Here, $f$, $g_{j}$ are quadratic functions of $\Phi^{\prime}$ and $a$. $n_{\text{cons}}$ is the total number of constraints. $\Phi^{\prime}\in\mathbb{R}^{n_{\Phi^{\prime}}}$ is the discretized current potential, and $a\in\mathbb{R}^{n_{a}},b_{j}\in\mathbb{R}$ are auxiliary variables for constructing complex problems. $c_{j}$ are constant weights. Unlike in NESCOIL or REGCOIL, $f$ or $g_{j}$ may be convex or non-convex. This allows QUADCOIL to target a greater range of objectives and constraints.
@@
With proper choice of $f,g_{j},b_{j}$, and $a$, (8) can represent a wide range of coil optimization problems. These include the REGCOIL problem (6), the L-1-regularized problem (7), and interesting new problems, such as minimizing the stored magnetic energy $E(\Phi^{\prime})$ under a local error constraint:
@@ 2.3 Shor relaxation
Shor relaxation comes with one drawback. The number of free variables in the relaxed CP scales as $\mathcal{O}[(n_{\Phi^{\prime}}+n_{a})^{2}+n_{\text{cons}}]$. As a result, compared to existing current potential methods, QUADCOIL scales poorly with high $n_{\Phi^{\prime}}$ and $n_{a}$. This may become a problem in future works applying finite element basis to QUADCOIL. See Fig 2 for the solve time of (20) in QUADCOIL as a function of $n$ and $n_{\text{cons}}$. ($m=0$ in this problem).
@@ Appendix C Relaxation procedures
Following this procedure, we can rewrite $f,g_{j}$ from the QCQP in Eq. (8) into $\text{tr}(FX),\text{tr}(G_{j}X)$, and transform the equation into a linear program (LP) on a closed, non-convex, compact domain $\mathbb{D}$:
$\min_{X\in\mathbb{D},b_{j}\in\mathbb{R}}\left[\text{tr}(FX)+\sum_{j}^{n_{\text{cons}}}c_{j}b_{j}\right].$ (48)
@@
$\begin{split}\mathbb{D}_{p}=\{X|X\in\mathbb{R}^{n_{x}^{2}},\text{tr}(G_{j}X)+d_{j}b_{j}\leq 0,\}\end{split}$ (50)
@@
$\min_{X\in\text{conv}(\mathbb{D}),b_{j}\in\mathbb{R}}\left[\text{tr}(FX)+\sum_{j}^{n_{\text{cons}}}c_{j}b_{j}\right].$ (52)
@@
$\begin{split}\min&\left[\text{tr}(FX)+\sum_{j}^{n_{\text{cons}}}c_{j}b_{j}\right],\\ X&\in\mathbb{D}_{p}\cap\text{conv}(\mathbb{D}_{n}),\\ b_{j}&\in\mathbb{R}.\end{split}$ (54)

### v6

@@ 2.2 QUADCOIL
$\begin{split}&\min_{\Phi^{\prime},a,b}\left[f(\Phi^{\prime},a)+c^{T}b\right]\\ &\text{subject to }\\ &g_{j}(\Phi^{\prime},a)+d_{j}^{T}b\leq e_{j},\\ &h_{k}(\Phi^{\prime},a)+m_{k}^{T}b=n_{k},\\ &\text{for all }j=1,...,n_{\text{ineq}},k=1,...,n_{\text{eq}}.\end{split}$ (8)
Here, $f$, $g_{j}$ are quadratic functions of $\Phi^{\prime}$ and $a$. $n_{\text{ineq}}$ and $n_{\text{eq}}$ are the total number of inequality and equality constraints. $\Phi^{\prime}\in\mathbb{R}^{n_{\Phi^{\prime}}}$ is the discretized current potential, and $a\in\mathbb{R}^{n_{a}},b\in\mathbb{R}^{n_{b}}$ are auxiliary variables for constructing complex problems. $c,d_{j}$ and $e_{j}$ are constants. Unlike in NESCOIL or REGCOIL, $f$ or $g_{j}$ may be convex or non-convex. This allows QUADCOIL to target a greater range of objectives and constraints.
@@
With proper choice of $f,g_{j},b$, and $a$, (8) can represent a wide range of coil optimization problems. These include the REGCOIL problem (6), the L-1-regularized problem (7), and interesting new problems, such as minimizing the stored magnetic energy $E(\Phi^{\prime})$ under a local error constraint:
@@ 2.3 Shor relaxation
Shor relaxation comes with one drawback. The number of free variables in the relaxed CP scales as $\mathcal{O}[(n_{\Phi^{\prime}}+n_{a})^{2}+n_{b}]$. As a result, compared to existing current potential methods, QUADCOIL scales poorly with high $n_{\Phi^{\prime}}$ and $n_{a}$. This may become a problem in future works applying finite element basis to QUADCOIL. See Fig 2 for the solve time of (20) in QUADCOIL as a function of $n$. ($m=0$ in this problem).
@@ Appendix C Relaxation procedures
Following this procedure, we can rewrite $f,g_{j},h_{k}$ from the QCQP in Eq. (8) into $\text{tr}(FX),\text{tr}(G_{j}X),\text{tr}(H_{k}X)$, and transform the equation into a linear program (LP) on a closed, non-convex, compact domain $\mathbb{D}$:
$\min_{X\in\mathbb{D},b\in\mathbb{R}^{n_{b}}}\left[\text{tr}(FX)+c^{T}b\right].$ (48)
@@
$\begin{split}\mathbb{D}_{p}=\{X|X\in\mathbb{R}^{n_{x}^{2}},&\text{tr}(G_{j}X)+d_{j}^{T}b\leq e_{j},\\ &\text{tr}(H_{k}X)+m_{k}^{T}b=n_{k},\}\end{split}$ (50)
@@
$\min_{X\in\text{conv}(\mathbb{D}),b\in\mathbb{R}^{n_{b}}}\left[\text{tr}(FX)+c^{T}b\right].$ (52)
@@
$\begin{split}\min&\left[\text{tr}(FX)+c^{T}b\right],\\ X&\in\mathbb{D}_{p}\cap\text{conv}(\mathbb{D}_{n}),\\ b&\in\mathbb{R}^{n_{b}}.\end{split}$ (54)